# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ III-A Achievable rate and SINR regions in the absence of power constraints
Consider the model and notation in Example 3. To simplify notation, we drop the arguments $({\bm{v}}_{n})_{n\in\mathcal{N}}$ because the sets $(\mathcal{Y}_{n}=\{{\bm{v}}_{n}\})_{n\in\mathcal{N}}$ are assumed to be singletons. For instance, ${\bm{c}}_{n}({\bm{v}}_{n})$, $b_{n}({\bm{v}}_{n})$, and $\sigma_{n}({\bm{v}}_{n})$ for each $n\in\mathcal{N}$ is simply denoted by ${\bm{c}}_{n}$, $b_{n}$, and $\sigma_{n}$, respectively. As a result, by defining $(b_{1},\ldots,b_{N})=:\bm{b}\in{\mathbb{R}}_{++}^{N}$, $(\bm{c}_{1},\ldots,\bm{c}_{N})=:\bm{C}\in{\mathbb{R}}_{+}^{N\times N}$, and $(\sigma_{1},\ldots,\sigma_{N})=:\bm{\sigma}\in{\mathbb{R}}_{++}^{N}$, the mapping $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}^{N}$ in the fixed point equation ${\bm{p}}={\mathrm{diag}(\bm{s})}T({\bm{p}})$ in (19) simplifies to $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}^{N}_{++}:{\bm{p}}\mapsto{\bm{Mp}}+{\bm{u}},$ where ${\bm{M}}:=\mathrm{diag}(\bm{b})^{-1}\bm{C}^{t}$ is referred to as the interference matrix, and $\bm{u}:=\mathrm{diag}({\bm{b}})^{-1}{\bm{\sigma}}.$ For this mapping $T$, we readily verify that its corresponding asymptotic mapping $T_{\infty}$ (see Definition 2) is given by $T_{\infty}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}^{N}:{\bm{p}}\mapsto{\bm{Mp}},$ and thus the achievable SINR region $\mathcal{S}$ in (22) simplifies to
@@
which is the characterization of the achievable SINR region discussed in [13, Ch. 5] for another particular case of (15). Therefore, (22) can be seen as a generalization of the results in [13, Ch. 5] to possibly nonlinear systems. Our generalization is able to cover, for example, optimal beamforming strategies, such as those studied in [21].
For the affine interference model in the previous example, we can also obtain an easily verifiable, sufficient condition for convexity of $\mathcal{S}$ in (23), which is a simple consequence of [28, Theorem 4.3] (Fact 5 in the Supplemental Material) – see also Example 1:
@@
If the interference matrix $\bm{M}$} in Example 4 is an inverse Z-matrix, then, by [28, Theorem 4.3] (Fact 5 in the Supplemental Material), the function $\mathcal{D}_{+}^{N\times{\mathbb{N}}}\to{\mathbb{R}}_{+}:{\bm{D}}\mapsto\rho({\bm{DM}})$ is convex. Therefore, the SINR region $\mathcal{S}$ in (23) is convex because this set is the intersection of the level set of a convex function with the convex cone ${\mathbb{R}}_{++}^{N}$.
If we use the standard nonlinear bijection $\mathrm{rate}=\log(1+\mathrm{SINR})$ (NOTE: here $\log$ represents the natural logarithm, so the rate is expressed in nats per symbol), in which case we say that we are treating interference as noise (single-user decoder), we can express the relationship between rates and SINR of all users as ${\bm{D}}=e^{\bm{R}}-{\bm{I}}$, where ${\bm{D}}=\mathrm{diag}(s_{1},\ldots,s_{N})$ is the diagonal matrix of SINR values, ${\bm{R}}=\mathrm{diag}(r_{1},\ldots,r_{N})$ is the corresponding diagonal matrix of rates, and $\bm{I}$} denotes the identity matrix. Therefore, recalling (22), we can express the set of achievable rates $\mathcal{R}$ as the following set
@@
The results in [29, Lemmas 3.2 and 3.3] (Fact 7 in the Supplemental Material) imply that, for any two general interference mappings $G_{1}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ and $G_{2}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ such that $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~G_{1}({\bm{x}})\leq G_{2}({\bm{x}})$, we have $\rho(G_{1})\leq\rho(G_{2})$. Therefore, $h$ is monotone in the sense that $(\forall{\bm{x}}\in{\mathbb{R}}^{N})(\forall{\bm{y}}\in{\mathbb{R}}^{N})~{\bm{x}}\leq{\bm{y}}\Rightarrow h({\bm{x}})\leq h({\bm{y}})$. To prove that $h$ is also convex, choose two vectors $({\bm{x}}_{1},{\bm{x}}_{2})\in{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}$ and a scalar $\alpha\in~]0,1[$ arbitrarily. We establish convexity of $h$ as follows:
@@
where the first inequality is a consequence of monotonicity of $h$, and the second inequality follows from the assumption of convexity of ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}T_{\infty})$ on ${\mathbb{R}}_{+}^{N}$. Therefore, we deduce from [40, p. 86][41, Theorem 2.1.3(vi)] that the function $h\circ g:{\mathbb{R}}^{N}\to{\mathbb{R}}_{+}$, where
@@
is also convex because $h\circ g$ is the composition of the monotonic convex function $h$ and the function $g$, which is convex in each coordinate. Convexity of $h\circ g$ implies convexity of the function $f$ because $f$ is the restriction of the convex function $h\circ g$ to the nonnegative orthant. ∎
The discussion above establishes that convexity of the mapping ${\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}:{\bm{s}}\mapsto\rho(\mathrm{diag}({\bm{s}})T_{\infty})$ implies the convexity of the set $\mathcal{S}$ in (22) because $\mathcal{S}$ is the level set of a convex function. As an immediate implication of this result and Proposition 4, we deduce that the set $\mathcal{R}$ in (24) is also convex. Moreover, Remark 2 provides a simple criterion for establishing convexity in specific cases. In more detail, for affine interference models such as those described in Example 4, the sets $\mathcal{S}$ and $\mathcal{R}$ are convex whenever the interference matrix $\bm{M}$} is an inverse Z-matrix. For later reference, we formalize these results below.
@@
If the asymptotic mapping $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ is such that the function ${\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}:{\bm{s}}\mapsto\rho(\mathrm{diag}({\bm{s}})T_{\infty})$ is convex, then the sets $\mathcal{S}$ and $\mathcal{R}$ in (22) and (24), respectively, are convex. In particular, for the interference model in Example 4, the sets $\mathcal{S}$ and $\mathcal{R}$ are convex if the interference matrix $\bm{M}$} in the definition of the linear asymptotic mapping $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto{\bm{Mx}}$ is an inverse Z-matrix.
III-B Achievable rate and SINR regions with power constraints
In the previous section, we discussed existing results proving that there exists at most one power vector ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$ able to provide a given SINR requirement ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$, and, if it exists, this power vector is precisely the fixed point of the mapping ${\mathrm{diag}(\bm{s})}T$ – see (21). Therefore, to describe the set $\mathcal{S}_{\mathcal{P}}$ of positive achievable SINR values with power constraints, we have to remove from the set $\mathcal{S}$ in (22) all points $\bm{s}$} corresponding to feasible SINR levels obtained with a power vector ${\bm{p}}\notin\mathcal{P}$, where $\mathcal{P}$ is defined in (18). More precisely, verifying whether $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$ is nonempty is insufficient. If $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$, we also need to determine whether the unique fixed point ${\bm{p}}^{\star}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$, which is the only power allocation able to provide the SINR requirement represented by the vector $\bm{s}$}, satisfies ${\bm{p}}^{\star}\in\mathcal{P}$. From [23, Proposition 5] (Fact 10 in the Supplemental Material) we know that ${\bm{p}}^{\star}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$ and ${\bm{p}}\in\mathcal{P}$ if and only if the conditional eigenvalue $\lambda$ associated with ${\mathrm{diag}(\bm{s})}T$ and $\|\cdot\|$ satisfies $\lambda\leq 1$, and Proposition 1 shows that $\lambda\leq 1$ if and only if $\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})\leq 1$, where $T_{\|\cdot\|}$} is the general interference mapping defined in (20). As a result, the set of positive SINR values achievable under the power constraint $\mathcal{P}$, excluding the points on the boundary ${\mathbb{R}}_{+}^{N}\backslash{\mathbb{R}}^{N}_{++}$ of the nonnegative cone ${\mathbb{R}}_{+}^{N}$, can be expressed in the following compact form:
@@ III-B Achievable rate and SINR regions with power constraints
and thus ${\bm{0}}<{\bm{x}}<\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}}({\bm{x}})$. This last inequality proves that there exists a scalar $\lambda>1$ satisfying $\lambda{\bm{x}}\leq\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}}({\bm{x}})$, which implies $\rho(\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}})>1$ as a consequence of [29, Lemmas 3.2 and 3.3] (Fact 7(iii) in the Supplemental Material). Therefore, we have ${\bm{s}}+{\bm{\epsilon}}\notin\mathcal{S}_{\mathcal{P}}$ for every ${\bm{\epsilon}}\in{\mathbb{R}}_{++}^{N}$, so $\bm{s}$} lies on the weak Pareto boundary of $\mathcal{S}_{\mathcal{P}}$. ∎
If $T_{\|\cdot\|}$} is norm inducing in the sense of Definition 5, with induced norm denoted by $\|\cdot\|_{T_{\|\cdot\|}}$, which satisfies $(\forall{\bm{s}}\in{\mathbb{R}}_{+}^{N})~\|{\bm{s}}\|_{T_{\|\cdot\|}}=\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ as shown in Proposition 3, the closure $\overline{{\mathcal{S}}_{\mathcal{P}}}$ of the set $\mathcal{S}_{\mathcal{P}}$ is given by
@@
Assume that $T_{\|\cdot\|}$} in (20) is norm inducing. Then the set $\overline{\mathcal{R}_{\mathcal{P}}}$ of achievable rates is a compact convex set.
@@
Owing to space constraints, we direct readers to the Supplementary Material for the proof, which follows a similar approach to that of Proposition 4. ∎
@@
With $K=N$ and ${\bm{a}}_{1},\ldots,{\bm{a}}_{K}$ as the standard basis vectors, Corollary 2 could be partially deduced from the findings in [9]. Even in this particular case of the results we present, our approach offers some significant advantages. First, convexity of the SINR and rate regions can be identified even in scenarios with strong interference and high SNR, which is a regime that cannot be covered by results in previous studies. For example, the hypothetical interference matrix ${\bm{M}}=\left[\begin{matrix}2&10\\ 10^{-1}&1\end{matrix}\right]$ has a component (in the first row, second column) corresponding to strong interference compared to all other components of the matrix. Nevertheless, ${\bm{M}}+{\bm{ua}}_{n}^{t}$ is an inverse Z-matrix for every $n\in\{1,\ldots,K\}$ if the components of all vectors $({\bm{a}}_{n})_{n\in\{1,\ldots,N\}}$ are sufficiently small, which can be guaranteed by making the transmit power sufficiently large. Therefore, we are neither in the low SNR regime nor in the low interference regime, yet the conditions of Corollary 2 holds. Second, our approach focuses on the spectral radius of a single nonlinear mapping (see (26)), which not only streamlines the notation but also accommodates more flexible power constraints. More generally, the characterization of the feasible rate and SINR regions through the spectral radius of the nonlinear mappings in (26) and (27) does not require affine interference models or convex polytopes as the power constraints, so our characterization is inaccessible to previous results that do not use the concept of nonlinear spectral radius.

### v3

@@ III-B Achievable rate and SINR regions in the absence of power constraints
Consider the model and notation in Example 6. To simplify notation, we drop the arguments $({\bm{v}}_{n})_{n\in\mathcal{N}}$ because the sets $(\mathcal{Y}_{n}=\{{\bm{v}}_{n}\})_{n\in\mathcal{N}}$ are assumed to be singletons. For instance, ${\bm{c}}_{n}({\bm{v}}_{n})$, $b_{n}({\bm{v}}_{n})$, and $\sigma_{n}({\bm{v}}_{n})$ for each $n\in\mathcal{N}$ is simply denoted by ${\bm{c}}_{n}$, $b_{n}$, and $\sigma_{n}$, respectively. As a result, by defining $(b_{1},\ldots,b_{N})=:\bm{b}\in{\mathbb{R}}_{++}^{N}$, $(\bm{c}_{1},\ldots,\bm{c}_{N})=:\bm{C}\in{\mathbb{R}}_{+}^{N\times N}$, and $(\sigma_{1},\ldots,\sigma_{N})=:\bm{\sigma}\in{\mathbb{R}}_{++}^{N}$, the mapping $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}^{N}$ in the fixed point equation ${\bm{p}}^{\star}={\mathrm{diag}(\bm{s})}T({\bm{p}}^{\star})$ in (24) simplifies to $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}^{N}_{++}:{\bm{p}}\mapsto{\bm{Mp}}+{\bm{u}},$ where ${\bm{M}}:=\mathrm{diag}(\bm{b})^{-1}\bm{C}^{t}$ is referred to as the interference matrix, and $\bm{u}:=\mathrm{diag}({\bm{b}})^{-1}{\bm{\sigma}}.$ For this mapping $T$, we readily verify that its corresponding asymptotic mapping $T_{\infty}$ (see Definition 2) is given by $T_{\infty}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}^{N}:{\bm{p}}\mapsto{\bm{Mp}},$ and thus the achievable SINR region $\mathcal{S}$ in (27) simplifies to
@@
which is the characterization of the achievable SINR region discussed in [35, Ch. 5] for another particular case of (20). Therefore, the characterization of the set $\mathcal{S}$ in (27) – which does not assume linear asymptotic mappings $T_{\infty}$ and includes the linear mapping $T_{\infty}$ in this Example as a special case – can be seen as a generalization of the results in [35, Ch. 5] to possibly nonlinear systems. In particular, (27) covers the model in Example 7, and it can also accommodate optimal beamforming strategies, such as those studied in [27].
For the affine interference model in the previous example, we can also obtain an easily verifiable, sufficient condition for convexity of $\mathcal{S}$ in (28): [^5] The results in Remark 2 may be useful for establishing convexity for interference models beyond those considered in this study.
@@
If the interference matrix $\bm{M}$} in Example 8 is an inverse Z-matrix, then, by Fact 5 in the Supplemental Material, the function $\mathcal{D}_{+}^{N\times{\mathbb{N}}}\to{\mathbb{R}}_{+}:{\bm{D}}\mapsto\rho({\bm{DM}})$ is convex. Therefore, the SINR region $\mathcal{S}$ in (28) is convex because this set is the intersection of the level set of a convex function with the convex cone ${\mathbb{R}}_{++}^{N}$.
More generally, if we use the standard nonlinear bijection $\mathrm{rate}=\log(1+\mathrm{SINR})$ (NOTE: here $\log$ represents the natural logarithm, so the rate is expressed in nats per symbol), in which case we say that we are treating interference as noise (single-user decoder), we can express the relationship between rates and SINR of all users as ${\bm{D}}=e^{\bm{R}}-{\bm{I}}$, where ${\bm{D}}=\mathrm{diag}(s_{1},\ldots,s_{N})$ is the diagonal matrix of SINR values, ${\bm{R}}=\mathrm{diag}(r_{1},\ldots,r_{N})$ is the corresponding diagonal matrix of rates, and $\bm{I}$} denotes the identity matrix. Therefore, recalling (27), we can express the set of achievable rates $\mathcal{R}$ as the following set
@@
The results in Fact 7 in the Supplemental Material imply that, for any two general interference mappings $G_{1}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ and $G_{2}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ such that $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~G_{1}({\bm{x}})\leq G_{2}({\bm{x}})$, we have $\rho(G_{1})\leq\rho(G_{2})$. Therefore, $h$ is monotone in the sense that $(\forall{\bm{x}}\in{\mathbb{R}}^{N})(\forall{\bm{y}}\in{\mathbb{R}}^{N})~{\bm{x}}\leq{\bm{y}}\Rightarrow h({\bm{x}})\leq h({\bm{y}})$. To prove that $h$ is also convex, choose two vectors $({\bm{x}}_{1},{\bm{x}}_{2})\in{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}$ and a scalar $\alpha\in~]0,1[$ arbitrarily. We establish convexity of $h$ as follows:
@@
where the first inequality is a consequence of monotonicity of $h$, and the second inequality follows from the assumption of convexity of ${\bm{s}}\mapsto\rho({\mathrm{diag}(\bm{s})}T_{\infty})$ on ${\mathbb{R}}_{+}^{N}$. Therefore, we deduce from [6, p. 86][43, Theorem 2.1.3(vi)] that the function $h\circ g:{\mathbb{R}}^{N}\to{\mathbb{R}}_{+}$, where
@@
is also convex because $h\circ g$ is the composition of the monotonic convex function $h$ and the mapping $g$, which is convex in each coordinate. Convexity of $h\circ g$ implies convexity of the function $f$ because $f$ is the restriction of the convex function $h\circ g$ to the nonnegative orthant. ∎
The discussion above establishes that convexity of the mapping ${\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}:{\bm{s}}\mapsto\rho(\mathrm{diag}({\bm{s}})T_{\infty})$ implies the convexity of the set $\mathcal{S}$ in (27) because $\mathcal{S}$ is the level set of a convex function. As an immediate implication of this result and Proposition 4, we deduce that the set $\mathcal{R}$ in (29) is also convex. Moreover, Remark 4 provides a simple criterion for establishing convexity in special cases. In more detail, for affine interference models such as those described in Example 8, the sets $\mathcal{S}$ and $\mathcal{R}$ are convex whenever the interference matrix $\bm{M}$} is an inverse Z-matrix. For later reference, we formalize these results below.
@@
If the asymptotic mapping $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ is such that the function ${\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}:{\bm{s}}\mapsto\rho(\mathrm{diag}({\bm{s}})T_{\infty})$ is convex, then the sets $\mathcal{S}$ and $\mathcal{R}$ in (27) and (29), respectively, are convex. In particular, for the interference model in Example 8, the sets $\mathcal{S}$ and $\mathcal{R}$ are convex if the interference matrix $\bm{M}$} in the definition of the linear asymptotic mapping $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{p}}\mapsto{\bm{Mp}}$ is an inverse Z-matrix.
III-C Achievable rate and SINR regions with power constraints
In the previous section, we discussed existing results proving that there exists at most one power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$ able to provide a given SINR requirement ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$, and, if it exists, this power vector is precisely the fixed point of the mapping ${\mathrm{diag}(\bm{s})}T$ – see (26). Therefore, to describe the set $\mathcal{S}_{\mathcal{P}}$ of positive achievable SINR values with power constraints, we have to remove from the set $\mathcal{S}$ in (27) all points $\bm{s}$} corresponding to feasible SINR levels obtained with a power vector ${\bm{p}}^{\star}\notin\mathcal{P}$, where $\mathcal{P}$ is defined in (23). More precisely, verifying whether $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$ is nonempty is insufficient. If $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$, we also need to determine whether the unique fixed point ${\bm{p}}^{\star}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$, which is the only power allocation able to provide the SINR requirement represented by the vector $\bm{s}$}, satisfies ${\bm{p}}^{\star}\in\mathcal{P}$. From Fact 10 in the Supplemental Material we know that ${\bm{p}}^{\star}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$ and ${\bm{p}}^{\star}\in\mathcal{P}$ if and only if the conditional eigenvalue $\lambda$ associated with ${\mathrm{diag}(\bm{s})}T$ and $\|\cdot\|$ satisfies $\lambda\leq 1$, and Proposition 1 shows that $\lambda\leq 1$ if and only if $\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})\leq 1$, where $T_{\|\cdot\|}$} is the general interference mapping defined in (25). As a result, the set of positive SINR values achievable under the power constraint $\mathcal{P}$, excluding the points on the boundary ${\mathbb{R}}_{+}^{N}\backslash{\mathbb{R}}^{N}_{++}$ of the nonnegative cone ${\mathbb{R}}_{+}^{N}$, can be expressed in the following compact form:
@@ III-C Achievable rate and SINR regions with power constraints
and thus ${\bm{0}}<{\bm{x}}<\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}}({\bm{x}})$. This last inequality proves that there exists a scalar $\lambda>1$ satisfying $\lambda{\bm{x}}\leq\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}}({\bm{x}})$, which implies $\rho(\mathrm{diag}({\bm{s}}+{\bm{\epsilon}}){T_{\|\cdot\|}})>1$ as a consequence of Fact 7(iii) in the Supplemental Material. Therefore, we have ${\bm{s}}+{\bm{\epsilon}}\notin\mathcal{S}_{\mathcal{P}}$ for every ${\bm{\epsilon}}\in{\mathbb{R}}_{++}^{N}$, so $\bm{s}$} lies on the weak Pareto boundary of $\mathcal{S}_{\mathcal{P}}$. ∎
If $T_{\|\cdot\|}$} is norm-inducing in the sense of Definition 5, with induced norm denoted by $\|\cdot\|_{T_{\|\cdot\|}}$, which satisfies $(\forall{\bm{s}}\in{\mathbb{R}}_{+}^{N})~\|{\bm{s}}\|_{T_{\|\cdot\|}}=\rho({\mathrm{diag}(\bm{s})}{T_{\|\cdot\|}})$ as shown in Proposition 3, the closure $\overline{{\mathcal{S}}_{\mathcal{P}}}$ of the set $\mathcal{S}_{\mathcal{P}}$ is given by
@@
Assume that $T_{\|\cdot\|}$} in (25) is norm-inducing. Then the set $\overline{\mathcal{R}_{\mathcal{P}}}$ of achievable rates is a compact convex set.
@@
Now, follow the exact same steps below (30) to conclude that the set $\overline{\mathcal{R}_{\mathcal{P}}}$ is convex. Boundedness of ${\bm{r}}^{\star}\in\overline{\mathcal{R}_{\mathcal{P}}}$ follows immediately from boundedness of closed balls $\{{\bm{x}}\in{\mathbb{R}}^{N}\mid\|{\bm{x}}\|_{T_{\|\cdot\|}}\leq 1\}$, nonnegativity of vectors in $\overline{\mathcal{R}_{\mathcal{P}}}$, and the fact that $x\to\infty$ implies $e^{x}\to\infty$. Therefore, being closed and bounded in the standard (finite-dimensional) Euclidean space, $\overline{\mathcal{R}_{\mathcal{P}}}$ is compact in this space. ∎
@@
With $K=N$ and ${\bm{a}}_{1},\ldots,{\bm{a}}_{K}$ as the standard basis vectors, Corollary 2 could be partially deduced from the findings in [16]. Even in this particular case of the results we present, our approach offers some significant advantages. First, convexity of the SINR and rate regions can be identified even in scenarios with strong interference and high SNR, which is a regime that cannot be covered by results in previous studies. For example, the hypothetical interference matrix ${\bm{M}}=\left[\begin{matrix}2&10\\ 10^{-1}&1\end{matrix}\right]$ has a component (in the first row, second column) corresponding to strong interference compared to all other components of the matrix. Nevertheless, ${\bm{M}}+{\bm{ua}}_{n}^{t}$ is an inverse Z-matrix for every $n\in\{1,\ldots,K\}$ if the components of all vectors $({\bm{a}}_{n})_{n\in\{1,\ldots,N\}}$ are sufficiently small, which can be guaranteed by making the transmit power sufficiently large. Therefore, we are neither in the low SNR regime nor in the low interference regime, yet the conditions of Corollary 2 hold. Second, our approach focuses on the spectral radius of a single nonlinear mapping (see (31)), which not only streamlines the notation but also accommodates more flexible power constraints. More generally, the characterization of the feasible rate and SINR regions through the spectral radius of the nonlinear mappings in (31) and (32) does not require affine interference models or convex polytopes as the power constraints, so our characterization is inaccessible to previous results that do not use the concept of nonlinear spectral radius.